# Lifelong Learning of Video Diffusion Models From a Single Video Stream

### v2

@@ 7 Conclusion
This paper explores lifelong learning of diffusion models from video streams, which has never been investigated to the best of our knowledge. We establish the feasibility of learning video diffusion models from a single autocorrelated video stream in a lifelong fashion. Our most successful learning paradigm is remarkably simple and uses no specialty techniques beyond a minimal replay buffer. To promote further research into this area, we introduce three datasets that test how different video stream characteristics affect model and lifelong learning algorithm performance.
The ability to effectively train video diffusion models from a continuous data stream holds significant promise, particularly for world model learning [27, 42, 30] and compute-efficient foundation model adaptation [69]. Our findings suggest that learning generative models from strongly correlated data streams may not be as difficult as previously assumed, at least for certain data streams and model classes, echoing results from Bornschein et al. [12] and Carreira et al. [18] on discriminative model lifelong learning. We expect sample quality to improve significantly with data and compute scale and also with more advanced model architectures, optimization techniques, and continual learning algorithms. Further analysis of lifelong learned video diffusion model behavior, compute-efficient video model adaptation methods, and lifelong learning video models on more complex video streams are promising avenues of future inquiry.
Acknowledgments
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), the Canada CIFAR AI Chairs Program, Inverted AI, MITACS, the Department of Energy through Lawrence Berkeley National Laboratory, and Google. This research was enabled in part by technical support and computational resources provided by the Digital Research Alliance of Canada Compute Canada (alliancecan.ca), the Advanced Research Computing at the University of British Columbia (arc.ubc.ca), and Amazon.
Supplementary Material
@@ Appendix A The ColorKL Metric
$\frac{1}{|\mathcal{C}|}\sum_{c_{old}\in\mathcal{C}}D_{\mathrm{KL}}\left(p_{*}(c_{new}|c_{old})\;\|\ p_{model}(c_{new}|c_{old})\right)$
@@ Appendix B Additional Experiment Details
We employ the U-Net video diffusion architecture from Harvey et al. [35], which in turn is based on the U-Net image diffusion architecture of Nichol and Dhariwal [56]. We train the model using the epsilon prediction objective [39] and generate video samples during test time using Euler’s method sampler of Karras et al. [44].
@@
As computing all metrics on the entirety of the video streams is prohibitively expensive, we select 1,000 video subsequences from the evaluation stream and calculate the metrics on those video frames for all reported metrics. All training and sampling seeds compute the metrics on the same set of video subsequences. For datasets without significant changes in video frame details throughout the video stream (Lifelong Bouncing Balls (O), Lifelong 3D Maze), we select the first 1,000 video subsequences from the evaluation video stream. For datasets with significant changes in video frame details throughout the video stream (Lifelong Bouncing Balls (C), Lifelong PLAICraft), we evenly select 1,000 video subsequences across the entire evaluation video stream with equal spacing.
@@ Appendix C Additional Dataset Curation Details
The train and test set of Lifelong Bouncing Balls is generated by randomly sampling the two balls’ initial positions and velocities and deterministically updating them to satisfy the conservation of momentum.
The Lifelong 3D Maze dataset was created by concatenating two 10-hour-long Lifelong 3D Maze YouTube videos [26, 65] at the point where a maze from the first video is solved and the maze from the second video begins, as the maze screensaver teleports the agent to a newly generated maze on the completion of the previous maze. This concatenation ensures that there is no sudden switch in the environment dynamics (a slight perceptual switch exists as shown in Fig. 7).
@@
The Lifelong PLAICraft dataset was constructed by concatenating multiple consecutive gameplay session recordings from the same player of the PLAICraft server [36]. All game settings, such as viewing distance, shader configurations, and recording parameters (e.g., resolution and aspect ratio), were consistent across sessions. The dataset includes recordings of two players, Alex and Kyrie, who played in the same survival world alongside hundreds of other players. In their initial gameplay sessions, their starting locations were randomly assigned but could later be altered through travel or teleportation. As in a typical Minecraft survival world, the environment had no boundaries; landscapes were procedurally generated and remained unchanged once created. Player states, including inventories and spawn locations, were preserved between consecutive sessions.
@@ Appendix D Additional Quantitative Results
Method FVD Loss minADE ColorKL FVD Loss minADE ColorKL
Offline Learning 5.5 ±0.2 6.2e-5 ±2e-6 1.63 ±0.03 0.005 ±0.001 5.4 ±0.1 6.2e-5 ±2e-6 1.58 ±0.04 0.005 ±0.0
Experience Replay 5.0 ±0.1 6.3e-5 ±1e-6 1.71 ±0.05 0.004 ±0.0 5.0 ±0.2 6.2e-5 ±1e-6 1.68 ±0.01 0.004 ±0.0
No Replay 5.3 ±0.2 6.7e-5 ±1e-6 1.79 ±0.01 0.005 ±0.0 5.4 ±0.1 6.6e-5 ±1e-6 1.75 ±0.02 0.004 ±0.001
Full Replay 5.3 ±0.2 6.2e-5 ±0.0 1.68 ±0.01 0.004 ±0.0 5.1 ±0.2 6.2e-5 ±0.0 1.62 ±0.01 0.004 ±0.001
Table 5: Lifelong Bouncing Balls (O) performance metrics. The left and right columns respectively denote training and test video stream results computed across two training and three sampling random seeds. Lower is better for all metrics.
@@
Method FVD Loss minADE ColorKL FVD Loss minADE ColorKL
Offline Learning 6.9 ±0.3 6.5e-5 ±2e-6 1.92 ±0.09 0.007 ±0.002 8.0 ±0.5 6.5e-5 ±2e-6 2.01 ±0.09 0.006 ±0.001
Experience Replay 6.1 ±0.3 7.4e-5 ±0.0 1.89 ±0.03 0.005 ±0.0 6.7 ±0.3 7.5e-5 ±1e-6 1.96 ±0.01 0.005 ±0.001
No Replay 343.5 ±1.6 2.3e-2 ±1e-3 2.59 ±0.09 0.022 ±0.0 341.1 ±1.8 2.2e-2 ±1e-3 2.68 ±0.1 0.023 ±0.001
Full Replay 5.6 ±0.3 7.9e-5 ±2e-6 1.99 ±0.01 0.005 ±0.001 6.3 ±0.4 7.9e-5 ±2e-6 2.03 ±0.0 0.006 ±0.002
Table 6: Lifelong Bouncing Balls (C) performance metrics. The left and right columns respectively denote training and test video stream results computed across two training and three sampling random seeds. Lower is better for all metrics.
This section presents the train and test video stream performances of offline training (Offline Learning), lifelong learning using experience replay with limited replay buffer size (Experience Replay) [^2] This configuration is referred to as Lifelong Learning in the main text’s Sec. 5., lifelong learning using experience replay with unlimited replay buffer size (Full Replay), and lifelong learning without the use of a replay buffer (No Replay) on all datasets. Additional qualitative results appear in Appendix G.
D.1 Lifelong Bouncing Balls
For Lifelong Bouncing Balls (O), all learning algorithms attain comparable results. However, for Lifelong Bouncing Balls (C), No Replay performs significantly worse on all metrics than the rest as No Replay’s generated ball colors are different from those in the conditioning frames when the conditioning frames contain ball colors observed early in the video stream (refer to Fig. 12 and Fig. 16). This suggests that when lifelong learning on video streams that contain non-repeating details, mechanisms that preserve past knowledge are necessary regardless of how simple the video streams might be. This conclusion is further supported by Fig. 9(b), where the fully trained No Replay model performs progressively worse on data from earlier in the training stream.
@@ D.2 Lifelong 3D Maze
We observe that all learning algorithms attain comparable results on the test stream. However, No Replay performs significantly worse on the train stream metrics. This is in part due to the way the Lifelong 3D Maze dataset is constructed. The dataset is built by concatenating two 10-hour-long YouTube videos, and the test set only comprises unobserved frames from the second video. While the videos are nearly identical, there are very subtle perceptual differences (refer to Fig. 7). Because the training stream contains frames from the first video and the test stream contains frames from the second video, No Replay performs better on the test stream videos than on the train stream videos as it overfits to the perceptual details of the second video.
D.3 Lifelong PLAICraft
@@
Method FVD KVD Loss FVD KVD Loss
Offline Learning 93.5 ±1.3 66.6 ±0.9 0.0056 ±0.0 89.2 ±1.0 62.7 ±0.8 0.0056 ±0.0
Experience Replay 98.0 ±3.2 62.8 ±2.7 0.0068 ±0.0001 98.8 ±2.4 62.3 ±1.5 0.0059 ±0.0
No Replay 220.0 ±0.9 168.2 ±0.8 0.0086 ±0.0 90.3 ±1.0 54.6 ±0.6 0.0058 ±0.0
Full Replay 97.4 ±1.3 67.5 ±1.2 0.0055 ±0.0 92.3 ±2.1 63.2 ±1.6 0.0057 ±0.0
Table 7: Lifelong 3D Maze performance metrics. Results are computed across two training and three sampling random seeds. Lower is better for all metrics.
@@
Method FVD KVD Loss FVD KVD Loss
Offline Learning 123.2 ±1.1 10.6 ±0.3 0.0330 191.8 ±1.6 14.9 ±0.3 0.0403
Experience Replay 107.3 ±1.2 8.3 ±0.2 0.0334 170.7 ±1.4 12.0 ±0.6 0.0408
No Replay 400.7 ±3.1 27.3 ±0.3 0.0487 485.2 ±3.7 42.8 ±0.7 0.0530
Full Replay 121.8 ±1.8 10.0 ±0.3 0.0329 185.0 ±0.5 13.0 ±0.1 0.0398
Table 8: Lifelong PLAICraft performance metrics. Results are computed across one training and three sampling seeds. Lower is better for all metrics.
We observe that Offline Learning and experience replay-based methods perform similarly but No Replay performs significantly worse. Fig. 9(d) shows the final models’ performances on video frames at different training stream indices. To further analyze model behaviors, we plot the fully trained models’ performance metrics on video frames from different parts of the training stream. Interestingly, the relative performance difference between different training stream frame indices for the lifelong learned models, even No Replay, mirrors that of the offline-learned model. This suggests that video stream underfitting, not forgetting, is the primary challenge in effective learning of Lifelong PLAICraft as we can assume that the offline-learned model does not suffer from forgetting.
Appendix E Training Time Performance Metrics

### v3

@@ 7 Conclusion
This paper explores lifelong learning of diffusion models from continuous video streams, which has never been investigated to the best of our knowledge. We establish the feasibility of learning video diffusion models from a single autocorrelated video stream in a lifelong fashion. Our lifelong learning approach is simple, using no specialty techniques beyond a minimal replay buffer, and is robust across various data streams, model architectures, parameter sizes, and optimizer configurations. To promote further research into this area, we introduce four single video datasets that test how different video stream characteristics affect lifelong learning of video diffusion models.
The ability to train video diffusion models directly from continuous data streams holds significant promise, particularly for video world model learning in embodied agents [1, 3, 26, 42, 29] and compute-efficient foundation model adaptation [71]. Our results suggest that learning generative models from strongly correlated data streams may be less challenging than previously assumed, echoing prior results on discriminative lifelong learning [13, 18, 34]. This paper’s preliminary investigations lay the groundwork for advancing lifelong learning of video generative models by massively scaling continuous video streams and model sizes, incorporating advanced diffusion architectures, and developing novel learning algorithms.
@@ Appendix A The ColorKL Metric
$\frac{1}{|\mathcal{C}|}\sum_{c_{old}\in\mathcal{C}}D_{\mathrm{KL}}\left(p_{*}(c_{new}|c_{old})\;\|\ p_{model}(c_{new}|c_{old})\right)$ (3)
@@ Appendix B Additional Experiment Details
As computing all metrics on the entirety of the video streams is prohibitively expensive, we select 1,000 video subsequences from the train and test streams and calculate the metrics on those video frames for all reported metrics. All training and sampling seeds compute the metrics on the same set of video subsequences. For datasets without significant changes in video frame details throughout the video stream (Lifelong Bouncing Balls (O)), we select the first 1,000 video subsequences from the evaluation video stream. For datasets with significant changes in video frame details throughout the video stream (Lifelong Bouncing Balls (C), Lifelong Drive, Lifelong PLAICraft), we evenly select 1,000 video subsequences across the entire evaluation video stream with equal spacing. For Lifelong 3D Maze, the train stream evaluation is done across 1,000 video subsequences evenly selected from the stream and the test stream evaluation is done across the first 1,000 video subsequences from the stream due to the subtle train stream perceptual changes detailed in Appendix C.
All models are optimized with AdamW with learning rate 1e-4. The optimizers for U-Net models have weight decay of 1e-5, whereas the optimizers for Transformer models have weight decay of 0 for Lifelong Bouncing Balls and 1e-6 for all other experiments. All U-Net models have the num_res_blocks hyperparameter set to 1. The num_channels hyperparameter is set to 64, 192, 128, 128, 128 for Lifelong Bouncing Balls (O), Lifelong Bouncing Balls (C), Lifelong 3D Maze, Lifelong Drive, and Lifelong PLAICraft experiments respectively. All VDT models have the num_layers hyperparameter set to 12. The hidden_layer_size and patch_size hyperparameters are set to 384, 384, 640, 1024, 1024 and 2, 2, 4, 4, 8 for Lifelong Bouncing Balls (O), Lifelong Bouncing Balls (C), Lifelong 3D Maze, Lifelong Drive, and Lifelong PLAICraft experiments respectively. We applied gradient clipping with a threshold of 1 to lifelong learned VDT models for Lifelong PLAICraft and Lifelong Drive experiments to mitigate infrequent but sporadic gradient norm spikes.
The U-Net and VDT experiments for Lifelong Bouncing Balls took 3 days to complete on a single GeForce RTX 2080 GPU. The U-Net and VDT experiments for Lifelong 3D Maze took 7 days to complete on two Tesla V100 GPUs. The U-Net experiments for Lifelong Drive and Lifelong PLAICraft took 4 days to complete on 4 A5000 GPUs, and the VDT experiments for Lifelong Drive and Lifelong PLAICraft took 6 days to complete on 4 L40 GPUs. There was no substantial runtime difference between Offline Learning and Lifelong Learning since both methods employ the same batch sizes and the same number of gradient steps.
@@ Appendix C Additional Dataset Curation Details
Lifelong Bouncing Balls
videos were created by randomly sampling the two balls’ initial positions and velocities and deterministically updating them to satisfy the conservation of momentum.
Lifelong 3D Maze
was created by concatenating two 10-hour-long Lifelong 3D Maze YouTube videos [25, 67] at the point where a maze from the first video is solved and the maze from the second video begins, as the maze screensaver teleports the agent to a newly generated maze on the completion of the previous maze. This concatenation ensures that there is no sudden switch in the environment dynamics (a slight perceptual switch exists as shown in fig. 4).
@@
Lifelong Drive
was constructed by first downloading a YouTube video that features a vehicle driving on a series of highways from Chongqing to Shanghai [79], then cropping and downsampling the video such that its frame dimensions became 512x512. The drive is nearly continuous, with subtle and rare discontinuities occurring when the driver takes breaks in between very long drives. In these scenarios, the video either resumes at the same spot where the vehicle was parked or at a nearby highway location.
Lifelong PLAICraft
was constructed by concatenating multiple consecutive gameplay session recordings from the same player of the PLAICraft server [36]. All game settings, such as viewing distance, shader configurations, and recording parameters (e.g., resolution and aspect ratio), were consistent across sessions. The dataset includes recordings of two players, Alex and Kyrie, who played in the same survival world alongside hundreds of other players. In their initial gameplay sessions, their starting locations were randomly assigned but could later be altered through travel or teleportation. As in a typical Minecraft survival world, the environment had no boundaries; landscapes were procedurally generated and remained unchanged once created. Player states, including inventories and spawn locations, were preserved between consecutive sessions.
@@ Appendix D Additional Quantitative Results
This section presents the train and test video stream performances of offline training (Offline Learning), lifelong learning using experience replay with limited buffer size (Experience Replay) [^2] This configuration is referred to as Lifelong Learning in the main text’s section 5., lifelong learning using experience replay with unlimited buffer size (Full Replay), and lifelong learning without the use of a replay buffer (No Replay) for U-Net models on all datasets.
We make two high-level observations. Firstly, No Replay cannot effectively learn from datasets that are not largely stationary due to forgetting, even on simple datasets like Lifelong Bouncing Balls (C). This suggests that when lifelong learning on video streams that contain non-repeating details, mechanisms that preserve past knowledge are necessary, regardless of how simple the video streams might be. Secondly, we find that Experience Replay and Full Replay performances are not substantially different across the benchmarks, suggesting that storing 5 to 20 percent of video stream frames can be sufficient to have a performant model when lifelong learning video diffusion models.
D.1 Lifelong Bouncing Balls
@@
Method FVD Loss×${}^{\times\!10^{-\!5}}$ minADE ColorKL FVD Loss×${}^{\times\!10^{-\!5}}$ minADE ColorKL
Offline Learning 4.5 ±0.1 6.3 ±0.1 1.74 ±0.03 0.006 ±0.001 4.7 ±0.2 6.2 ±0.2 1.71 ±0.06 0.006 ±0.001
No Replay 4.7 ±0.1 6.7 ±0.1 1.91 ±0.01 0.006 ±0.0 5.0 ±0.2 6.7 ±0.1 1.84 ±0.03 0.005 ±0.0
Experience Replay 4.9 ±0.2 6.3 ±0.2 1.82 ±0.03 0.005 ±0.000 4.7 ±0.2 6.2 ±0.3 1.81 ±0.03 0.005 ±0.000
Full Replay 4.7 ±0.1 6.3 ±0.1 1.80 ±0.01 0.004 ±0.001 4.6 ±0.1 6.2 ±0.0 1.76 ±0.02 0.003 ±0.001
Table 6: Lifelong Bouncing Balls (O) performance metrics for U-Net models. The left and right columns respectively denote training and test video stream results computed across two training and three sampling random seeds.
@@
Method FVD Loss×${}^{\times\!10^{-\!5}}$ minADE ColorKL FVD Loss×${}^{\times\!10^{-\!5}}$ minADE ColorKL
Offline Learning 5.8 ±0.3 6.5 ±0.1 2.04 ±0.09 0.007 ±0.002 5.9 ±0.2 6.5 ±0.1 2.14 ±0.10 0.007 ±0.001
No Replay 357.4 ±1.8 2240 ±110 2.61 ±0.08 0.021 ±0.002 343.6 ±1.2 2252 ±108 2.73 ±0.11 0.022 ±0.0
Experience Replay 5.0 ±0.1 7.4 ±0.1 2.03 ±0.00 0.005 ±0.001 5.7 ±0.2 7.5 ±0.1 2.06 ±0.00 0.005 ±0.000
Full Replay 5.0 ±0.1 8.0 ±0.1 2.08 ±0.03 0.005 ±0.001 5.6 ±0.2 7.9 ±0.2 2.12 ±0.0 0.006 ±0.001
Table 7: Lifelong Bouncing Balls (C) performance metrics for U-Net models. The left and right columns respectively denote training and test video stream results computed across two training and three sampling random seeds.
@@ D.2 Lifelong 3D Maze
Method FVD JEDi Loss FVD JEDi Loss
Offline Learning 34.2 ±1.3 0.083 ±0.002 0.005 ±0.0 28.5 ±1.2 0.060 ±0.001 0.005 ±0.0
No Replay 130.6 ±0.6 0.142 ±0.002 0.009 ±0.0 36.9 ±1.9 0.104 ±0.004 0.006 ±0.0
Experience Replay 41.2 ±0.4 0.073 ±0.006 0.006 ±0.0 30.7 ±0.1 0.061 ±0.003 0.006 ±0.0
Full Replay 34.0 ±0.3 0.080 ±0.002 0.006 ±0.0 35.2 ±0.5 0.075 ±0.006 0.006 ±0.0
Table 8: Lifelong 3D Maze metrics for experience replay with different replay buffer sizes for U-Net models. The metrics are computed across one training and three sampling seeds.
D.3 Lifelong Drive
@@
Method FVD JEDi Loss FVD JEDi Loss
Offline Learning 15.3 ±0.2 0.071 ±0.000 0.029 ±0.0 17.9 ±0.3 0.102 ±0.001 0.026 ±0.0
No Replay 322.8 ±1.2 0.608 ±0.002 0.041 ±0.0 183.4 ±1.7 0.343 ±0.000 0.029 ±0.0
Experience Replay 17.0 ±0.2 0.072 ±0.000 0.030 ±0.0 21.8 ±0.3 0.106 ±0.000 0.027 ±0.0
Full Replay 15.8 ±0.0 0.072 ±0.000 0.030 ±0.0 19.3 ±0.2 0.096 ±0.000 0.026 ±0.0
Table 9: Lifelong Drive metrics for experience replay with different replay buffer sizes for U-Net models. The metrics are computed across one training and three sampling seeds.
D.4 Lifelong PLAICraft
Train Stream Test Stream
Method FVD JEDi Loss FVD JEDi Loss
Offline Learning 59.7 ±0.2 0.874 ±0.007 0.034 ±0.0 119.8 ±1.3 1.119 ±0.004 0.042 ±0.0
No Replay 239.5 ±1.5 2.038 ±0.014 0.047 ±0.0 270.6 ±1.9 2.281 ±0.004 0.050 ±0.0
Experience Replay 62.9 ±0.6 0.876 ±0.003 0.034 ±0.0 130.8 ±1.1 1.189 ±0.003 0.042 ±0.0
Full Replay 67.8 ±0.7 0.867 ±0.004 0.034 ±0.0 131.6 ±0.7 1.170 ±0.001 0.042 ±0.0
Table 10: Lifelong PLAICraft metrics for experience replay with different replay buffer sizes for U-Net models. The metrics are computed across one training and three sampling seeds.
Appendix E Orthogonal Replay
We denote the AdamW-optimized experience replay objective as Experience Replay, and denote the Orthogonal-AdamW optimized experience replay objective as Orthogonal Replay.
Train Stream Test Stream
Method FVD Loss×${}^{\times\!10^{-\!5}}$ minADE ColorKL FVD Loss×${}^{\times\!10^{-\!5}}$ minADE ColorKL
Experience Replay (U) 4.9 ±0.2 6.3 ±0.2 1.82 ±0.03 0.005 ±0.000 4.7 ±0.2 6.2 ±0.3 1.81 ±0.03 0.005 ±0.000
Orthogonal Replay (U) 4.8 ±0.1 6.4 ±0.0 1.82 ±0.02 0.006 ±0.000 4.9 ±0.3 6.2 ±0.0 1.75 ±0.00 0.005 ±0.000
Experience Replay (T) 5.5 ±0.1 6.6 ±0.2 2.22 ±0.05 0.005 ±0.001 5.0 ±0.3 6.5 ±0.2 2.16 ±0.06 0.005 ±0.001
Orthogonal Replay (T) 5.5 ±0.2 6.4 ±0.2 2.17 ±0.02 0.006 ±0.002 5.3 ±0.1 6.3 ±0.2 2.11 ±0.01 0.007 ±0.001
Table 11: Lifelong Bouncing Balls (O) metrics computed across two training and three sampling seeds. The configurations (U) and (T) denote U-Net and Transformer-based diffusion architecture.
Train Stream Test Stream
Method FVD Loss×${}^{\times\!10^{-\!5}}$ minADE ColorKL FVD Loss×${}^{\times\!10^{-\!5}}$ minADE ColorKL
Experience Replay (U) 5.0 ±0.1 7.4 ±0.1 2.03 ±0.00 0.005 ±0.001 5.7 ±0.2 7.5 ±0.1 2.06 ±0.00 0.005 ±0.000
Orthogonal Replay (U) 5.2 ±0.1 8.2 ±0.1 2.10 ±0.11 0.004 ±0.001 6.0 ±0.1 8.2 ±0.1 2.12 ±0.07 0.004 ±0.000
Experience Replay (T) 5.3 ±0.1 8.1 ±0.0 2.29 ±0.03 0.006 ±0.001 6.4 ±0.3 8.1 ±0.0 2.27 ±0.03 0.006 ±0.000
Orthogonal Replay (T) 5.4 ±0.2 7.7 ±0.1 2.23 ±0.01 0.006 ±0.001 6.0 ±0.2 7.8 ±0.1 2.25 ±0.00 0.006 ±0.001
Table 12: Lifelong Bouncing Balls (C) metrics computed across two training and three sampling seeds. The configurations (U) and (T) denote U-Net and Transformer-based diffusion architecture.
Train Stream Test Stream
Method FVD JEDi Loss FVD JEDi Loss
Experience Replay (U-Net) 41.2 ±0.4 0.073 ±0.006 0.006 ±0.0 30.7 ±0.1 0.061 ±0.003 0.006 ±0.0
Orthogonal Replay (U-Net) 41.3 ±0.4 0.080 ±0.003 0.006 ±0.0 31.8 ±0.7 0.084 ±0.001 0.006 ±0.0
Experience Replay (Transformer) 32.1 ±0.9 0.082 ±0.005 0.006 ±0.0 30.2 ±0.8 0.070 ±0.001 0.006 ±0.0
Orthogonal Replay (Transformer) 34.5 ±0.6 0.067 ±0.003 0.006 ±0.0 29.7 ±1.2 0.071 ±0.002 0.006 ±0.0
Table 13: Lifelong 3D Maze metrics computed across one training and three sampling seeds.
Train Stream Test Stream
Method FVD JEDi Loss FVD JEDi Loss
Experience Replay (U-Net) 17.0 ±0.2 0.072 ±0.000 0.030 ±0.0 21.8 ±0.3 0.106 ±0.000 0.027 ±0.0
Orthogonal Replay (U-Net) 18.8 ±0.1 0.078 ±0.000 0.030 ±0.0 25.6 ±0.2 0.125 ±0.000 0.026 ±0.0
Experience Replay (Transformer) 10.7 ±0.1 0.023 ±0.000 0.029 ±0.0 12.9 ±0.1 0.039 ±0.000 0.025 ±0.0
Orthogonal Replay (Transformer) 11.1 ±0.1 0.027 ±0.000 0.029 ±0.0 13.6 ±0.2 0.044 ±0.000 0.025 ±0.0
Table 14: Lifelong Drive metrics computed across one training and three sampling seeds.
Train Stream Test Stream
Method FVD JEDi Loss FVD JEDi Loss
Experience Replay (U-Net) 62.9 ±0.6 0.876 ±0.003 0.034 ±0.0 130.8 ±1.1 1.189 ±0.003 0.042 ±0.0
Orthogonal Replay (U-Net) 69.2 ±1.1 0.842 ±0.006 0.035 ±0.0 136.3 ±0.9 1.206 ±0.001 0.043 ±0.0
Experience Replay (Transformer) 21.9 ±0.4 0.197 ±0.003 0.025 ±0.0 26.3 ±0.3 0.208 ±0.001 0.024 ±0.0
Orthogonal Replay (Transformer) 21.9 ±0.1 0.182 ±0.002 0.024 ±0.0 26.0 ±0.3 0.192 ±0.001 0.024 ±0.0
Table 15: Lifelong PLAICraft metrics computed across one training and three sampling seeds.
Appendix F Optimization Time Performance Metrics